OpenStudy (anonymous):

On his last three keyboarding tests Sean typed at 41 word per minute, 48 words per minute, and 50 words per minute. Which of these is a good estimate for how long it would take him to type a 460-word document? a.)about twleve minutes b.)about ten minutes c.)about five minutes d.)about 8 minutes

OpenStudy (anonymous):

i think you would first need to get an average because you want an estimate so you would add all 3 speeds and divide by 3 and then take that value divide it into 460.which would be about 10 minutes

OpenStudy (anonymous):

find the average words per minute, you can do that by adding up the three scores and dividing that by three. The once closest to the average is a good estimate. then to find for each time, take the average and multiply it by the number of minutes

OpenStudy (anonymous):

Divide 460 by the average typing rate. \[\frac{460}{\frac{1}{3} (41+48+50)} = 9.928.. \]